# M3D: Manifold-based Domain Adaptation with Dynamic Distribution for Non-Deep Transfer Learning in Cross-subject and Cross-session EEG-based Emotion Recognition

### v2

@@ Abstract
Emotion decoding using Electroencephalography (EEG)-based affective brain-computer interfaces (aBCIs) is crucial for affective computing but is hindered by EEG’s non-stationarity, individual variability, and the high cost of large-scale labeled data. Deep learning-based approaches, while effective, require substantial computational resources and large datasets, limiting their practicality. To address these challenges, we propose Manifold-based Domain Adaptation with Dynamic Distribution (M3D), a lightweight non-deep transfer learning framework. M3D includes four main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning. The data undergoes a transformation onto an optimal Grassmann manifold space, enabling dynamic alignment of the source and target domains. This process prioritizes both marginal and conditional distributions according to their significance, ensuring enhanced adaptation efficiency across various types of data. In the classifier learning, the principle of structural risk minimization is integrated to develop robust classification models. This is complemented by dynamic distribution alignment, which refines the classifier iteratively. Additionally, the ensemble learning module aggregates the classifiers obtained at different stages of the optimization process, which leverages the diversity of the classifiers to enhance the overall prediction accuracy. The proposed M3D framework is evaluated on two benchmark EEG emotion recognition datasets using two validation protocols (cross-subject single-session and cross-subject cross-session), as well as on a clinical EEG dataset of Major Depressive Disorder (MDD). Experimental results demonstrate that M3D outperforms traditional non-deep learning methods, achieving an average improvement of 4.47%, while achieving deep learning-comparable performance with significantly lower data and computational requirements. These findings highlight the potential of M3D to enhance the practicality and applicability of aBCIs in real-world scenarios.
Index Terms:
@@
11footnotetext:  Equal contributions.
I Introduction
Emotion, functioning as both a physiological and psychological state within an individual, serves as a rich repository of insights into mental and physical well-being [1, 2, 3]. Emotion recognition stands as a crucial method for objectively discerning human emotional states. Currently, the practice relies on the analysis of either non-physiological signals or physiological signals. In contrast to cues like facial expressions, voice intonation, and gestures, Electroencephalography (EEG) stands out as a non-invasive brain information collection technique [4]. Its strong correlation with human emotional states, coupled with its resistance to deception, positions EEG as a highly accurate means to objectively reflect human emotions [5, 6, 7].
@@ I Introduction
Early EEG-based emotion recognition methods mainly rely on conventional non-deep machine learning algorithms[11], including Support Vector Machine (SVM), K-Nearest Neighbor (KNN), and Linear Discriminant Analysis (LDA). In these studies, there is an underlying assumption that EEG samples are independent and identically distributed. However, variations among individuals significantly undermine the effectiveness of traditional machine learning methods in cross-subject emotion recognition tasks. This results in suboptimal performance and poor generalization effects. On the other hand, transfer learning techniques can be beneficial in transferring informative data from the source (training data) to the target domain (test data), which could reduce the variations between different distributions in the modeling. More related transfer learning methods are detailed in Section II.
Compared to deep learning-based transfer learning methods, non-deep learning-based transfer learning offers a more lightweight approach. Unlike deep learning methods, non-deep learning-based transfer requires less labeled data, which is better suited to the current conditions in EEG tasks. In this paper, we propose a novel non-deep learning-based transfer learning framework for cross-subject emotion recognition, named as Manifold-based Domain adaptation with Dynamic Distribution (M3D). The M3D model maps EEG data onto a manifold space, where both marginal and conditional distributions are dynamically analyzed and aligned. This process is coupled with classifier learning to minimize distributional differences between source and target domains. Subsequently, ensemble learning integrates the results to enhance the robustness of cross-subject and cross-session emotion recognition. Experimental results demonstrate that the proposed M3D model achieves performance comparable to existing deep learning-based methods. The main contributions of the paper are summarized below.
  • A novel non-deep learning-based transfer learning model ((M3D) is proposed to address the variations in the distribution of EEG signals across different sessions and subjects.
  • Both marginal distribution and conditional distribution are considered during domain alignment, where a dynamically adjusted weighting of the importance of the two distributions is incorporated.
  • An adaptive classifier learning together with an ensemble learning is introduced to iteratively refine the classifier learning process and leverage the strength of multiple learning models to optimize the prediction of data labels.
  • Extensive experiments are conducted on two well-known databases using two different validation protocols (cross-subject single-session and cross-subject cross-session), as well as a clinical EEG database of Major Depressive Disorder (MDD). The results obtained from the proposed M3D model (a non-deep learning approach) are comparable to those achieved by existing deep learning methods.
@@
Figure 1: The overall architecture of the proposed M3D model, which includes four mains modules. (1) Manifold Feature Transformation: incorporates TCA (Transfer Component Analysis) to obtain the embedded subspaces, and develops GFK (Geodesic Flow Kernel) algorithm to transform the original features into manifold features. (2) Dynamic Distribution Alignment: dynamically adjusts the importance of the marginal and conditional distributions by learning an adaptive factor $\mu$ during the alignment process. (3) Classifier Learning: summarizes structural risk minimization and distribution alignment to iteratively optimize the classifier. (4) Ensemble Learning: harmonizes the classification results from the iterative classifier learning process and generates a more reliable and robust classification performance. Here, $Source$ and $Target$ refer to the source and target data, respectively. $\mathbb{G}$ is the Grassmann manifold, projected by the geodesic function $\Psi(t)$ ($t\in[0,1]$). The points $\Psi(0)$ and $\Psi(1)$ correspond to the embedded representations of the source and target domains in $\mathbb{G}$. $x_{i,j}$ and $x_{k,l}$ are the samples from the source and target data,respectively. $z^{\infty}_{i,j,k,l}$ are the corresponding data representation in the manifold space by the geodesic transformation. $H^{\infty}_{K}$ denotes the reproducing kernel Hilbert space (RKHS), and $G$ is the computed geodesics kernel.
II RELATED WORK
@@
Domain adaptation stands out as the predominant method in transfer learning, frequently employed to address differences between the distributions of source and target data. It targets the reduction of distribution disparities between the two domains and align the feature distributions, achieving the assumption that the source and target data is independently and identically distributed [12]. Existing domain adaptation-based transfer learning algorithms can be generally categorized into two broad categories: deep transfer learning and non-deep transfer learning.
II-A Deep Transfer Learning Methods
@@
II-B Non-Deep Transfer Learning Methods
@@
Non-deep transfer learning methods align the feature distribution between source and target domains using conventional machine learning approaches. For example, Pan et al. [17] introduced a Transfer Component Analysis (TCA) algorithm, leveraging the reproducing kernel Hilbert space to learn transfer information. Through maximizing the Mean Discrepancy (MMD), it successfully mitigated the marginal distribution differences. Zheng et al. [18] introduced a Transductive Parameter Transfer (TPT) algorithm to build personalized classifiers for subjects from various source domains. This algorithm mapped the individualized classifier parameters to the target domain, minimizing the MMD between the source and target domains and adapting to marginal distribution. Wang et al. [19] introduced a Stratified Transfer Learning (STL) algorithm, which constructed a weak classifier with source data and making predictions on the target domain. Based on this foundation, it further conducted a more thorough exploration of intra-class relationships, leveraging class-relatedness to accomplish adaptive spatial dimensionality reduction.
Although the transfer learning strategies mentioned above effectively mitigate differences between the distributions of source and target data by considering marginal or conditional distributions, it’s crucial to note that the actual disparities in EEG between subjects primarily exist in their joint distribution [15]. Current domain adaptation algorithms either focus on marginal or conditional distributions, neglecting to assess the combined contributions in domain adaptation process. It would lead to inefficient data utilization and suboptimal results. To tackle this issue, He [15] proposed a Joint Distribution Adaptation (JDA), which involved matching these two distributions with equal weights to approximate the adjustment of the joint distribution [20]. Building upon the proposed JDA, other studies extended its capabilities by incorporating regularization [21], sparse representation [22], structural consistency [23], and domain-invariant clustering [24].
@@
However, many existing methods assume equal contributions from the conditional and marginal probability distributions, which is often unrealistic in real-world applications. The relative importance of these distributions in shaping the joint probability distribution varies based on domain similarity. For example, when the source and target domains are highly dissimilar, aligning the marginal distribution is more critical. Conversely, when the two domains exhibit high similarity, prioritizing conditional distribution alignment is more beneficial. Therefore, for better alignment between the two domains, dynamically adjusting the weighting of marginal and conditional distributions based on the actual data characteristics is a critical issue in non-deep transfer learning research.
@@
Moreover, most existing approaches attempt to align source and target distributions directly within the original feature space. However, this approach can be ineffective, as certain features may be distorted due to domain shifts or may be inherently domain-specific. To address this issue, several approaches leverage subspace geometrical structures, particularly the Grassmann manifold, for subspace learning. These methods aim to transform data into a new projection space where domain-invariant features can be extracted, minimizing distributional differences between source and target domains [25, 26, 27]. This highlights the necessity of feature transformation prior to distribution alignment.
@@
To address these critical challenges and enhance non-deep transfer learning performance in EEG-based emotion recognition, we propose the Manifold-based Domain adaptation with Dynamic Distribution (M3D) model. Based on the extracted differential entropy (DE) features from raw EEG signals, we first incorporate the TCA algorithm to reduce feature dimensionality and narrow the distribution gap between the features derived from the source and target domains. Subsequently, the features are embedded into the Grassmann manifold space through the computation of the Geodesic Flow Kernel (GFK)[28]. To enhance the adaptability of our framework, we introduce an adaptive factor that dynamically assesses the significance of both marginal and conditional distributions in the two domains. This factor aids in optimizing the embedding process and ensuring the effective representation of manifold features. Upon achieving a robust manifold feature representation, we proceed with classifier learning to optimize the classification performance under the complex conditions of cross-subject and cross-session scenarios. In this phase, conventional machine learning methods such as SVM, KNN, AdaBoost, Gaussian Naive Bayesian (GNB), Decision Tree (DT), and Bagging classifiers are employed as the initial classifiers. These classifiers are chosen due to their simplicity, computational efficiency, and strong adaptability to small datasets. Finally, the classification results derived from diverse parameter settings are harmonized through an ensemble learning approach. This ensemble strategy enhances the overall robustness and reliability of the proposed framework, ultimately contributing to more accurate and reliable classification outcomes.
III METHODOLOGY
@@
In this section, we will provide a detailed introduction to the proposed M3D model. As shown in Fig. 1, M3D includes four main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning. In the manifold feature transformation module, the input data initially undergoes a reduction to lower-dimensional features through TCA. Then, a manifold kernel $G$ is learned to facilitate the transformation of the reduced features into an optimal manifold space. In the dynamic distribution alignment module, Structural Risk Minimization (SRM) [29] is incorporated to dynamically align the feature distribution within the obtained manifold space. This adaptive alignment process enhances the model’s capacity to effectively capture and adapt to variations in the underlying data distribution. In the classifier learning module, a classifier is constructed using the aligned manifold features. Initially, conventional machine learning methods are employed, and the classifier is then further iteratively optimized by solving the given classification loss. In the ensemble learning module, all the optimized classification results from the iterative process in the classifier learning module are harmonized to generate a more robust and reliable performance. These are the final results for cross-subject and cross-session EEG-based emotion recognition.
III-A Manifold Feature Transformation
@@
In this paper, we introduce a manifold feature transformation process, as illustrated in Fig. 1. The source and target domains are represented as $D_{S}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ and $D_{T}=\{(x_{n+1},y_{n+1}),...,(x_{n+m},y_{n+m})\}$, respectively, with each domain featuring $D$-dimensional data representations. Through TCA[17], we map both domains into a lower $d$-dimensional feature space ($d\ll D$), denoted as $T_{S}$ and $T_{T}$. This process effectively reduces the dimensionality of the feature space, facilitating more efficient data analysis and interpretation. Subsequently, we embed the mapped $d$-dimensional feature subspace into the Grassmann manifold, denoted as $\mathbb{G}$. In this manifold space, our goal is to enhance the similarity between the source and target domains’ data as much as possible. Simultaneously, we aim to retain the unique attributes and inherent characteristics of each domain’s data. This approach ensures that, despite the alignment of the domains in a shared feature space, the distinctiveness and originality of each domain are preserved, maintaining the integrity and context of the data from both domains.
To make the data distributions of the two domains more similar in the Grassmann manifold, we either bring the source domain closer to the target domain or vice versa. Here, we construct geodesics within $\mathbb{G}$, treating the source and target domains as points in this manifold space. By navigating along these geodesics, we facilitate a systematic journey from the source domain towards the target domain, enabling a gradual harmonization of their data distributions within the Grassmann manifold and ensuring a smoother and more effective domain alignment. In the process of constructing geodesics, we apply the standard Euclidean metric to the Riemannian manifold. This allows the geodesic to be described by a parameterization function, $\Psi$, where $t$ ranges from 0 to 1, mapping to $\Psi(t)$ within the Grassmann manifold. Specifically, $\Psi(0)$ and $\Psi(1)$ correspond to the embedded representations of the source and target domains in $\mathbb{G}$, respectively. Thus, the geodesic delineates the trajectory by which $\Psi(0)$ evolves into $\Psi(1)$, representing the most direct path within the manifold space that connects these two domain representations. To ensure a smooth and continuous transformation across the manifold, the intermediate values of $t$ ($t\in(0,1)$), $\Psi(t)$ satisfies the following conditions:
@@
Here, $T_{T}\in\mathbb{R}^{M\times d}$ refer to the set of orthonormal base for the target domain. $\Gamma$ and $\Sigma$ are both $d\times d$ diagonal matrices, where their diagonal elements being $\cos{\theta_{k}}$ and $\sin{\theta_{k}}$ ($k=1,2,...,d$), respectively. The variable $\theta_{k}$ denotes the angles between the orthonormal bases $T_{S}$ and $T_{T}$, with the constraint $0\leq\theta_{1}\leq\theta_{2}\leq...\leq\theta_{d}\leq\dfrac{\pi}{2}$. $\theta_{k}$ serve as a metric for quantifying the alignment between the domains within the $\mathbb{G}$, indicating how closely the subspaces represented by $T_{S}$ and $T_{T}$ are oriented to one another.
The operation $\Psi(t)^{\top}x$ represents the projection of a feature vector $x$ from the $d$-dimensional feature space onto the Grassmann manifold space $\mathbb{G}$, where $x$ could be any data sample originating from either the source or the target domain. The selection of an appropriate value or range of values for $t$ becomes a significant challenge, as it directly influences the quality of the projection for $x$. The optimal choice of $t$ ensures that the projected data maintains a balance between adapting to the new domain and preserving its original domain characteristics. This balance is essential for effective domain adaptation, as it allows for the integration of domain-specific features while minimizing the loss of critical information during the transition.
Here, we introduce an integration strategy to identify the optimal $t$ or a set of $t$ points between any two given data samples $x_{i}$ and $x_{j}$, and utilize a geodesic kernel to facilitate the mapping process. This methodology effectively transforms the $d$-dimensional feature space into an infinite-dimensional feature space, which serves to mitigate domain drift (a common challenge in domain adaptation). For each pair of data samples, $x_{i}$ and $x_{j}$, we iteratively calculate their projections onto $\Psi(t)$, where $t$ varies incrementally from 0 to 1. Through this process, each data sample is mapped across a continuum within the Grassmann manifold space, capturing the evolution of its feature representation as it transitions from being aligned with the source domain ($t=0$) to being aligned with the target domain ($t=1$). The culmination of this iterative projection process results in the creation of two infinite-dimensional feature vectors, denoted as $z^{\infty}_{i}$ and $z^{\infty}_{j}$,
@@
where $z^{\infty}_{i}$ and $z^{\infty}_{j}$ encapsulate the comprehensive trajectory of each data sample across the manifold, providing a rich, multidimensional perspective on the domain adaptation process. This strategy enhances the adaptability and accuracy of domain adaptation by leveraging the inherent structure of the data within the manifold, thereby offering a more effective approach to overcoming the challenges of domain drift. From [28], the geodesic kernel is then determined as the inner product of the two infinite-dimensional vectors, as
@@
where $G\in\mathbb{R}^{D\times D}$ is a positive semi-definite matrix, computed as
@@
Here, $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are all diagonal matrices, and the corresponding diagonal elements are given as
@@
where $k$ is in the range of $[1,d]$. The computed geodesics kernel $G$ enables the effective transformation of a data sample $x$, initially positioned within the $d$-dimensional feature space, into the Grassmann manifold space, as
@@
III-B Dynamic Distribution Alignment
After the manifold feature transformation, the feature distributions between the source and target domains would exhibit greater similarity than seen with the original feature representation. Still, differences remain in both the marginal probability distribution and conditional probability distribution. Current distribution adaptation methods commonly assume equal importance for both the marginal distribution $P$ and conditional distribution $Q$ [32, 33]. Traditional distribution adaptation methods often treat the adaptation of marginal and conditional distributions as equally important, an assumption that may not always be applicable. In scenarios where significant differences exist between the source and target domains, the adaptation of the marginal distribution takes on increased importance. Conversely, in situations where the source and target domains are more closely aligned, the adaptation of the conditional distribution becomes more critical. Therefore, recognizing and adjusting the emphasis on marginal or conditional distribution adaptation based on the domains’ specific characteristics is essential for achieving successful distribution adaptation.

### v3

@@ Abstract
Emotion decoding using Electroencephalography (EEG)-based affective brain-computer interfaces (aBCIs) is crucial for affective computing but is hindered by EEG’s non-stationarity, individual variability, and the high cost of large-scale labeled data. Deep learning-based approaches, while effective, require substantial computational resources and large datasets, limiting their practicality. To address these challenges, we propose Manifold-based Domain Adaptation with Dynamic Distribution (M3D), a lightweight non-deep transfer learning framework. M3D includes four main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning. The data undergoes a transformation onto an optimal Grassmann manifold space, enabling dynamic alignment of the source and target domains. This process prioritizes both marginal and conditional distributions according to their significance, ensuring enhanced adaptation efficiency across various types of data. In the classifier learning, the principle of structural risk minimization is integrated to develop robust classification models. This is complemented by dynamic distribution alignment, which refines the classifier iteratively. Additionally, the ensemble learning module aggregates the classifiers obtained at different stages of the optimization process, which leverages the diversity of the classifiers to enhance the overall prediction accuracy. The proposed M3D framework is evaluated on three benchmark EEG emotion recognition datasets using two validation protocols (cross-subject single-session and cross-subject cross-session), as well as on a clinical EEG dataset of Major Depressive Disorder (MDD). Experimental results demonstrate that M3D outperforms traditional non-deep learning methods, achieving an average improvement of 6.67%, while achieving deep learning-comparable performance with significantly lower data and computational requirements. These findings highlight the potential of M3D to enhance the practicality and applicability of aBCIs in real-world scenarios.
11footnotetext:  Equal contributions.
{IEEEkeywords}
@@
1 Introduction
[…]
Emotion, functioning as both a physiological and psychological state within an individual, serves as a rich repository of insights into mental and physical well-being [1, 2, 3]. Emotion recognition stands as a crucial method for objectively discerning human emotional states. Currently, the practice relies on the analysis of either non-physiological signals or physiological signals. In contrast to cues like facial expressions, voice intonation, and gestures, Electroencephalography (EEG) stands out as a non-invasive brain information collection technique [4, 5]. Its strong correlation with human emotional states, coupled with its resistance to deception, positions EEG as a highly accurate means to objectively reflect human emotions [6, 7, 8].
@@ 1 Introduction
Figure 1: The overall architecture of the proposed M3D model, which includes four mains modules. (1) Manifold Feature Transformation, (2) Dynamic Distribution Alignment, (3) Classifier Learning, (4) Ensemble Learning. Here, $Source$ and $Target$ refer to the source and target data, respectively. $\mathbb{G}$ is the Grassmann manifold, projected by the geodesic function $\Psi(t)$ ($t\in[0,1]$). The points $\Psi(0)$ and $\Psi(1)$ correspond to the embedded representations of the source and target domains in $\mathbb{G}$. $x_{i,j}$ and $x_{k,l}$ are the samples from the source and target data,respectively. $z^{\infty}_{i,j,k,l}$ are the corresponding data representation in the manifold space by the geodesic transformation. $H^{\infty}_{K}$ denotes the reproducing kernel Hilbert space (RKHS), and $G$ is the computed geodesics kernel.
Early EEG-based emotion recognition methods mainly rely on conventional non-deep machine learning algorithms[12], including Support Vector Machine (SVM), K-Nearest Neighbor (KNN), and Linear Discriminant Analysis (LDA). In these studies, there is an underlying assumption that EEG samples are independent and identically distributed. However, variations among individuals significantly undermine the effectiveness of traditional machine learning methods in cross-subject emotion recognition tasks. This results in suboptimal performance and poor generalization effects. On the other hand, transfer learning techniques can be beneficial in transferring informative data from the source (training data) to the target domain (test data), which could reduce the variations between different distributions in the modeling. More related transfer learning methods are detailed in Section 2.
Compared to deep learning-based transfer learning methods, non-deep learning-based transfer learning offers a more lightweight approach. Unlike deep learning methods, non-deep learning-based transfer requires less labeled data, which is better suited to the current conditions in EEG tasks. In this paper, we propose a novel non-deep learning-based transfer learning framework for cross-subject emotion recognition, named as Manifold-based Domain adaptation with Dynamic Distribution (M3D). The M3D model maps EEG data onto a manifold space, where both marginal and conditional distributions are dynamically analyzed and aligned. This process is coupled with classifier learning to minimize distributional differences between source and target domains. Subsequently, ensemble learning integrates the results to enhance the robustness of cross-subject and cross-session emotion recognition. Experimental results demonstrate that the proposed M3D model achieves performance comparable to existing deep learning-based methods. The main contributions of the paper are summarized below. First, a novel non-deep learning-based transfer learning model ((M3D) is proposed to address the variations in the distribution of EEG signals across different sessions and subjects. Second, both marginal distribution and conditional distribution are considered during domain alignment, where a dynamically adjusted weighting of the importance of the two distributions is incorporated. Third, an adaptive classifier learning together with an ensemble learning is introduced to iteratively refine the classifier learning process and leverage the strength of multiple learning models to optimize the prediction of data labels. Last, extensive experiments are validated on three well-known databases, as well as a clinical EEG database, with an average accuracy improvement of 6.67%.
2 RELATED WORK
@@
Domain adaptation targets the reduction of distribution disparities between source and target domains, achieving the assumption that the source and target data is independently and identically distributed [13]. Existing methods can be categorized into deep transfer learning and non-deep transfer learning.
2.1 Deep Transfer Learning Methods
@@
2.2 Non-Deep Transfer Learning Methods
@@
Non-deep transfer learning methods align the feature distribution between source and target domains using conventional machine learning approaches. For example, Pan et al.[18] introduced a Transfer Component Analysis (TCA) algorithm to learn transfer information via reproducing kernel Hilbert space and mitigate the marginal distribution differences by maximizing the Mean Discrepancy (MMD). Zheng et al.[19] introduced a Transductive Parameter Transfer (TPT) algorithm, which maps personalized classifiers parameters to the target domain while minimizing the MMD for marginal distribution. Wang et al.[20] introduced a Stratified Transfer Learning (STL) algorithm to construct a weak classifier with source data and explore intra-class relationships for adaptive spatial dimensionality reduction.
@@
Although existing transfer learning strategies consider marginal or conditional distributions to reduce distribution differences, actual EEG disparities between subjects are in the joint distribution [16]. Current algorithms focus on either marginal or conditional distributions, leading to inefficient data use. He [16] proposed Joint Distribution Adaptation (JDA) to match marginal and conditional distributions equally for joint distribution adjustment [21]. However, existing methods assume equal contributions from conditional and marginal probability distributions, which is unrealistic. The importance of these distributions in the joint probability distribution varies based on domain similarity. Aligning marginal distribution is crucial when domains are highly dissimilar, while prioritizing conditional distribution is beneficial when domains are similar. Dynamically adjusting the weighting of marginal and conditional distributions based on data characteristics is a key issue in non-deep transfer learning.
@@
Most existing approaches align source and target distributions directly in the original feature space, which can be ineffective due to domain shifts. Some methods leverage the Grassmann manifold for subspace learning to transform data and extract domain-invariant features [22, 23, 24], highlighting the need for feature transformation before distribution alignment.
@@
To address these challenges and improve non-deep transfer learning for EEG-based emotion recognition, we propose the M3D model. It first uses TCA on differential entropy features from raw EEG signals to reduce dimensionality and narrow the distribution gap. Then, it embeds features into the Grassmann manifold via Geodesic Flow Kernel (GFK) [25]. An adaptive factor is introduced to assess the significance of marginal and conditional distributions for better embedding. Upon achieving a robust manifold feature representation, we proceed with classifier learning to optimize the classification performance under the complex conditions of cross-subject and cross-session scenarios. In this phase, conventional machine learning methods like SVM, KNN, etc., are used as initial classifiers. Finally, an ensemble learning approach harmonizes classification results, enhancing the robustness and reliability of the framework for more accurate outcomes.
3 METHODOLOGY
@@
The proposed M3D model, shown in Fig. 1, includes four main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning. First, input data is reduced to lower-dimensional features using TCA, and a manifold kernel $G$ is learned to map the features into an optimal manifold space. Then, dynamic distribution alignment uses Structural Risk Minimization (SRM) [26] to adaptively align feature distributions within this space, improving adaptability to data variations. Next, a classifier is built on the aligned features, initially trained with traditional methods and iteratively refined by optimizing classification loss. Finally, ensemble learning combines all optimized classifiers from the iterations to produce more robust results for cross-subject and cross-session EEG-based emotion recognition.
3.1 Manifold Feature Transformation
@@
In this paper, we introduce a manifold feature transformation process, as illustrated in Fig. 1. The source and target domains are represented as $D_{S}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ and $D_{T}=\{(x_{n+1},y_{n+1}),...,(x_{n+m},y_{n+m})\}$, respectively, with each domain featuring D-dimensional data representations.
Through TCA[18], we map both domains into a lower $d$-dimensional feature space ($d\ll D$), denoted as $T_{S}$ and $T_{T}$. This dimensionality reduction process simplifies data, facilitating more efficient data analysis and interpretation. Subsequently, we embed the $d$-dimensional feature subspace into the Grassmann manifold, denoted as $\mathbb{G}$. Our goal in this manifold space is two-fold. We aim to increase the similarity between the source and target domains’ data. Simultaneously, we want to retain the unique attributes and inherent characteristics of each domain’s data. This approach ensures that while aligning the domains in a shared feature space, their distinctiveness remains intact.
To make the data distributions of the two domains more similar in the Grassmann manifold, we construct geodesics within $\mathbb{G}$, treating the source and target domains as points in this manifold space. By moving along these geodesics, we can gradually align the data distributions of the two domains. When constructing geodesics, we apply the standard Euclidean metric to the Riemannian manifold. This allows us to describe the geodesic using a parameterization function $\Psi(t)$, where $t$ ranges from 0 to 1. Specifically, $\Psi(0)$ and $\Psi(1)$ correspond to the embedded representations of the source and target domains in $\mathbb{G}$, respectively. The geodesic is the path from $\Psi(0)$ to $\Psi(1)$, representing the most direct connection between the two domain representations in the manifold space. To ensure a smooth and continuous transformation across the manifold, the intermediate values of $t$ ($t\in(0,1)$), $\Psi(t)$ satisfies the following conditions:
@@
where $T_{T}\in\mathbb{R}^{M\times d}$ is the orthonormal base set for the target domain. $\Gamma$ and $\Sigma$ are $d\times d$ diagonal matrices. Their diagonal elements are $\cos{\theta_{k}}$ and $\sin{\theta_{k}}$ ($k=1,2,...,d$), respectively. The variable $\theta_{k}$ represents the angles between the orthonormal bases $T_{S}$ and $T_{T}$, with $0\leq\theta_{1}\leq\theta_{2}\leq...\leq\theta_{d}\leq\dfrac{\pi}{2}$. $\theta_{k}$ measures the alignment between the domains in $\mathbb{G}$.
The operation $\Psi(t)^{\top}x$ projects a feature vector $x$ from the $d$-dimensional feature space onto the Grassmann manifold space $\mathbb{G}$, where $x$ could be any data sample from the source or target domain. Selecting an appropriate $t$ value is crucial, as it directly influences the projection quality of $x$. The optimal $t$ balances domain adaptation and the preservation of original features. This balance is essential for effective domain adaptation, as it allows for the integration of domain-specific features while minimizing the loss of critical information during the transition.
We introduce an integration strategy to find the optimal $t$ or a set of $t$ points between any two data samples $x_{i}$ and $x_{j}$, and utilize a geodesic kernel to facilitate the mapping process. This method transforms the $d$-dimensional feature space into an infinite-dimensional feature space, reducing domain drift. For each pair of data samples $x_{i}$ and $x_{j}$, we iteratively calculate their projections onto $\Psi(t)$ as $t$ ranges from 0 to 1. Each data sample is mapped across the Grassmann manifold space, capturing the evolution of its feature representation during the domain transition. The result of this process is two infinite-dimensional feature vectors, denoted as $z^{\infty}_{i}$ and $z^{\infty}_{j}$,
@@
where $z^{\infty}_{i}$ and $z^{\infty}_{j}$ represent the entire trajectory of each data sample in the manifold, providing a comprehensive view of the domain adaptation process. This strategy improves the adaptability and accuracy of domain adaptation. From [25], the geodesic kernel is the inner product of the two infinite-dimensional vectors, as
@@
where $G\in\mathbb{R}^{D\times D}$ is a positive semi-definite matrix, calculated as
@@
Here, $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are diagonal matrices, and the corresponding diagonal elements are given as
@@
where $k\in[1,d]$. The computed geodesics kernel $G$ enables the effective transformation of a data sample $x$, initially positioned within the $d$-dimensional feature space, into the Grassmann manifold space, as
@@
3.2 Dynamic Distribution Alignment
After the manifold feature transformation, the feature distributions between the source and target domains would exhibit greater similarity than seen with the original feature representation. Still, differences remain in both the marginal probability distribution and conditional probability distribution. Current distribution adaptation methods commonly assume equal importance for both the marginal distribution $P$ and conditional distribution $Q$ [29, 30]. This assumption, however, may not always be applicable. In scenarios where significant differences exist between the source and target domains, the adaptation of the marginal distribution takes on increased importance. Conversely, in situations where the source and target domains are more closely aligned, the adaptation of the conditional distribution becomes more critical. Therefore, recognizing and adjusting the emphasis on marginal or conditional distribution adaptation based on the domains’ specific characteristics is essential for achieving successful distribution adaptation.